OpenStudy (anonymous):

the middle term of (k/2+2)8 is 1120 then value of k is

Parth (parthkohli):

Do you mean that the coefficient is 1120?

OpenStudy (nirmalnema):

(k/2+2)8=1120 (k+4)8=1120 4k+16=1120 4k=1120-16 4k=1104 k=1104/4 k=276